Mathematics 244 – Linear Algebra Definitions Quiz 8 – April 23, 2004

Your Name:

Do all your work on this sheet. You have 5 minutes.

Give the definition of the following term: x ∈ V is an eigenvector T : V → V with eigenvalue λ = 9.